1.A given aqueous solution has a density of x g/mL and y% solute by mass. Derive the mathematical equation relating the molarity of the solution, M, to the molality of the solution,m.

2. A solution prepared by adding a given mass of solute to 20 g of benzene freezes at temperature 0.3840C below the freezing point of pure benzene. The freezing point of solution prepared from the same mass of the solute and 20 grams of water freezes at -0.41850C. Assume that the solute is undissociated in benzene solution, but is completely dissociated into ions in water solutions. How many ions results from the ionization of one molecule in water solutions?


1
Expert's answer
2021-01-28T05:49:05-0500

Molarity is the number of moles of solute per liter of solution. Convert to density by multiplying the number of moles by the molecular mass of the compound. Convert density to molarity by converting to grams per liter and dividing by the molecular mass of the compound in grams.

Density = mass/volumep

"D = m\/V"

m = D*V


for a) m = 12*5.75 = 69.0 g


for b) 1 mL = 1 cm3 so 30.0 mL = 30 cm3.

Using that, m = 30*0.779 = 23.4 g